OpenStudy (sparklestaraa):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 |2x - 4| = 5 |3x - 4| = -12 Equation #1 and Equation #2 have the same number of solutions. Equation #1 has more solutions than equation #2. Equation #1 has fewer solutions than equation #2. None of the statements above describe the number of solutions to equations shown.

OpenStudy (danjs):

are you familiar with the absolute value ?

OpenStudy (sparklestaraa):

hmmm no not much but i think it is Equation #1 has more solutions than equation #2.

OpenStudy (danjs):

how did you arrive there

OpenStudy (sparklestaraa):

actually to be honest i just figured it was that answer because 3x-4 is larger than the other one so i thought maybe its reverse psychology

OpenStudy (danjs):

From the above, you see there are 2 solutions to the first...

OpenStudy (sparklestaraa):

always like no matter what the equation is

OpenStudy (danjs):

an absolute value, is like a magnitude of distance from the origin, you cant have a negative value for the magnitude

OpenStudy (danjs):

\[\left| -x \right| = x\]

OpenStudy (sparklestaraa):

ohhhh

OpenStudy (danjs):

now if there were a negative in front of the absolute value , it is possible \[-\left| -x \right| = -x\]

OpenStudy (danjs):

so how many solutions does equation 2 have?

OpenStudy (danjs):

\[\left| 3x-4 \right| = -12~~~~~~false\]

OpenStudy (sparklestaraa):

ohhhh ok :D